OpenStudy (logan13):

What is the solution to the equation x + 2.15 = –3.2?

OpenStudy (logan13):

x =

OpenStudy (logan13):

@inowalst

OpenStudy (inowalst):

Subtract -3.2-2.15

OpenStudy (logan13):

-1.13

OpenStudy (inowalst):

You sure about that?

OpenStudy (ilovemath14):

X+2.15=-3.2 -2.15 -2.15 ——————— X= -5.35

OpenStudy (inowalst):

^

OpenStudy (ilovemath14):

X= -5.35